Good evening, Please help me in these two problems
1) Mitch needs to section off his bean bushes from the rest of his garden. The bean section is 3 feet wide and half again as long. How much cord will he need?
2)The Smithson Building in the downtown area was resurfaced with 1,200 granite sections each 4' x 4’. Last year 30 sections needed to be repaired. What percent of the granite squares needed repair?
Faithfully yours
narah

Answer by tutorcecilia(2152)

1)Use the formula for the perimeter of a rectangle:
Perimeter = 2(length + width)
Width= 3 feet
Length = (3)(1/2)=1.5 feet
P=2(1.5+3)
P=2(4.5)
P=9.0ft

.
2)Use the percent formula:
Part/Whole=Percent/100
Plug-in the values:
30/1200=Percent (x)/100
Cross-multiply
(30)(100)=(1200)(x)
3000=1200x [Solve for x]
3000/1200=x
2.5%=x